OpenStudy (anonymous):

f (x) = 3x - 7 and g(x) = -2x -6. find (f o g)(4)

OpenStudy (anonymous):

please explain thank you

OpenStudy (anonymous):

f(g(x))

OpenStudy (anonymous):

g(x) = -2x-6 f(x), where x = g(x) 3(-2x-6)-7

OpenStudy (anonymous):

so when after I so that, I multiply it by 4

OpenStudy (anonymous):

let x = 4

OpenStudy (anonymous):

If I'm reading it right

OpenStudy (anonymous):

so it would be 3(-2(4) - 6) -7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Composition of functions: Write down the outside function, but where there is an x put ()'s. Inside the parenthesis write the inside function. For example: \[f(g(x))=3(-2x-6) -7=-6x-18-7=-6x-25\] Now substitute the 4 in for x

OpenStudy (anonymous):

ok so if I subsitute 4 into that equation \[f(g(x))=3(-2x-6) -7=-6x-18-7=-6x-25\] that would be my given answer?